# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

### v3

@@ 1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. (2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. (2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. (2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. (2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Zhang et al. (2023) introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture. Wei et al. (2023) explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al. (2022); Cong et al. (2021), their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al. (2021, 2022); Rampášek et al. (2022). Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks. As such, the research world has been actively looking for counterparts of Attention-based models.
@@
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012). State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
@@
STG learning can be treated as a complex process of understanding and forecasting the evolution of STG networks over time, which highly resembles the state space transition process where each state encapsulates the current configuration of a system and transitions represent changes over time. Deep learning-based SSSMs bring new horizons to STG learning tasks. However, great challenges pose in accurately and effectively adopting Mamba architectures for STG modeling, which is characterized as heterogeneous, dynamic, and non-stationary.
@@
Motivated by the excellent long-term contextual modeling ability and low computational overhead of SSSMs, we propose Spatial-Temporal Graph Mamba (STG-Mamba) as the first deep learning-based SSSM for effective data-centric STG learning. The main contributions of this work are summarized in the following:
@@
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend selective state space models to handle STG data. Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, with $N$ stacked Graph Selective State Space Block (GS3B) as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
  • A novel Spatial-Temporal Selective State Space Module (ST-S3M) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent adaptive spatial-temporal graph feature selection. ST-S3M is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
@@
  • We introduce KFGN as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context. In KFGN, the Dynamic Filter Linear serves as dynamic GNN module which generating input-specific graph structures. The KF-Upgrading mechanism models STG inputs from different temporal granularities as parallel streams, and the output embeddings are integrated through the statistical theory learning of Kalman Filtering for optimization.
@@ 2 Preliminaries
Spatial-Temporal Graph (STG). STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society. For a given graph network, if the node individuals have certain spatial connectivity at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network. Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network. Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network. $A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
Spatial-Temporal Graph System (STG system). We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System (STG system). In the theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response. For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
@@ 3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning. Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
3.1 Formulation of STG Selective State Space Models
@@
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Residual Encoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic Encoder module, and repeat it $N$ times. GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, State Space Selection Mechanism (SSSM), Element-Wise Concatenation, as well as th Residual Connections for every GS3B Encoder layer.
@@ 3.2 Kalman Filtering Graph Neural Networks
The motivation for employing Kalman Filtering-based methods in constructing GNNs stems from the need to enhance the reliability and accuracy of spatial-temporal forecasting. STG big data (such as traffic sensor records/weather station records/urban metro system In/Out records) often contains inherent biases and noises that existing methods typically overlook. By integrating Kalman Filtering-based optimization and upgrading, our KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances. This approach not only corrects dataset inherent errors but also significantly improves the model’s capacity to capture the complex inter-dependencies within STG patterns.
@@ 3.3 ST-S3M: Spatial-Temporal Selective State Space Module
ST-S3M plays the significant role of adaptive STG feature selection, serving as the counterpart of Attention mechanism. The architecture of ST-S3M is shown in Figure 1. We also give an vivid illustration of the STG feature selection procedure in Figure 4.
After being processed by KFGN module for dynamic spatiotemporal dependency modeling and statistical-based integration & upgrading. The generated embedding $\hat{h}_{fuse}$ is sent to ST-S3M for input-specific dynamic feature selection. In ST-S3M, the input STG embedding passes through a Linear Layer, followed by a splitting operation. Let $b$ denotes batch size, $l$ denotes sequence length, $d_{model}$ be the feature dimension, $d_{inner}$ be the model inner feature dimension, we have:
@@
$h_{\text{main}}^{{}^{\prime}}=\text{Conv1D}(h_{\text{main}})$
$h_{\text{main}}^{{}^{\prime\prime}}=\text{SiLU}(h_{\text{main}}^{{}^{\prime}})$ (15)
where $h_{\text{main}}^{{}^{\prime\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to state space selection mechanism for processing:
$h_{\text{sssm}}=\text{SSSM}(h_{main}^{{}^{\prime\prime}})$ (16)
where $h_{\text{sssm}}\in\mathbb{R}^{b\times l\times d_{inner}}$. On the same time, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$. The fused result is transformed through a Linear projection:
@@
The State Space Selection Mechanism in ST-S3M plays a pivotal role in adaptive spatiotemporal feature selection. We detail the parameter computation and updating process of the State Space Selection Mechanism in Algorithm 1. Further, the pseudo-code for the employed Selective Scan method is presented in Algorithm 2.
Algorithm 1 State Space Selection Mechanism Parameter Calculation and Upgrading.
Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $A_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$;
Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
1:  Compute input-independent params $A\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$:
2:  $A=-\exp(A_{log})$, $D=D$;
3:  Process input $x$ with Linear Layer $Input\_proj$:
5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,B,C$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $B,C\in\mathbb{R}^{b\times l\times n}$;
8:  Employ the $SelectiveScan$ function to proceed with $x,\Delta^{*},A,B,C,D$:
9:  $y=SelectiveScan(x,\Delta^{*},A,B,C,D)$;
10:  Return the Output $y$.
Algorithm 2 Selective Scan Algorithm
Input: $u,\Delta^{*},A,B,C,D$;
Output: $z$;
1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ and the second dim of $A$ as $n$;
2:  Discretize continuous parameters $A$ and $B$:
3:  $\Delta A\leftarrow\exp(\text{einsum}(\Delta^{*},A))$, where $\Delta A\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
4:  $\Delta B_{u}\leftarrow\text{einsum}(\Delta^{*},B,u)$, where $\Delta B_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
5:  Initialize state $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
6:  for $i$ from $1$ to $l$ do
7:     $x\leftarrow\Delta A[:,i]\times x+\Delta B_{u}[:,i]$;
8:     $z\leftarrow\text{einsum}(x,C[:,i,:])$;
9:     Append $z$ to the output list $z_{s}$;
10:  end for
11:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
12:  Adding the direct gain $D$:
13:  $z\leftarrow z+u\times D$
14:  Return $z$.
3.4 Computational Efficiency of STG-Mamba
@@ 4.1 Dataset Statistics and Baseline Methods
Datasets. For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al. (2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al. (2022), and weather station records across mainland China (KnowAir) Wang et al. (2020). Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system. The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval. HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval. KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018. Within the KnowAir dataset, there are three subsets. In this study, we exclusively use subset No. 3, covering the period from 1/Sep/2016 to 31/Jan/2017. The detailed dataset statistics and descriptions are summarized in Table 1.
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 48 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-4}$, with a weight decay of $1e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 100. The number of GS3B Encoder Layer is set as $N=4$. We use the historical ground-truth data from the past 12 steps (for historical recent/period/trend scales) to forecast the future 12 time steps. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.

### v4

@@ 1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the dynamic, heterogenous, and non-stationary nature of STG data, accurate and efficient STG forecasting has long been a challenging task.
@@
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Modern Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012). State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
@@
STG learning can be treated as a complex process of understanding and forecasting the evolution of STG networks over time, which highly resembles the state space transition process where each state encapsulates the current configuration of a system and transitions represent changes over time. Deep learning-based SSSMs bring new horizons to STG learning tasks. However, great challenges pose in accurately and effectively adopting SSSMs architectures for STG modeling.
@@
Motivated by the excellent long-term contextual modeling ability and low computational overhead of SSSMs, we propose Spatial-Temporal Graph Mamba (STG-Mamba) as the first deep learning-based SSSM for effective data-centric STG learning. The main contributions of this work are:
@@
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend SSSMs to handle STG data. Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, which takes $N$ stacked Graph Selective State Space Block (GS3B) as basic module, with the proposed Kalman Filtering Graph Neural Network (KFGN), the Spatial-Temporal Selective State Space Module (ST-S3M), and a simultaneous STG Feed-Forward connection to serialize and coordinate different internal modules.
  • A novel Spatial-Temporal Selective State Space Module (ST-S3M) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent adaptive spatial-temporal graph feature selection. The Graph Selective Scan algorithm within ST-S3M simultaneously receives graph information from KFGN through the Feed-Forward connection, the Feed-Forward graph information is then employed to assist more effective updating of the state transition matrix and control matrix.
@@
  • We introduce KFGN as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context. In KFGN, DynamicFilter-GNN serves as the module which generating input-specific dynamic graph structures. The KF-Upgrading mechanism models STG inputs from different temporal granularities as parallel streams, and the output embeddings are integrated through the statistical theory learning of Kalman Filtering for optimization.
@@ 2 Preliminaries
Spatial-Temporal Graph System (STG system). We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System. In theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response. For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
@@ 3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models in Section 3.1 and Appendix D as well. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning (Section 3.3). Finally, a detailed computational efficiency analysis of STG-Mamba in presented in Appendix B because of limited space.
3.1 Architecture of STG Selective State Space Models
@@
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Residual Encoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic Encoder module, and repeat it $N$ times. GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, Grapg State Space Selection Mechanism (GSSSM), Element-Wise Concatenation, Graph Information Feed-Forward, as well as the Residual Connection. The Graph Information Feed-Forward structure here is specifically designed for coordinating different modules’ information transmission and updating, ensuring the latest STG information is available for every module.
@@ 3.2 Kalman Filtering Graph Neural Networks
The motivation for employing Kalman Filtering-based methods in constructing GNNs stems from the need to enhance the reliability and accuracy of spatial-temporal forecasting. STG big data (i.e., traffic sensor records, weather station records, etc) often contains inherent biases and noises that existing methods typically overlook. By integrating Kalman Filtering-based optimization and upgrading, KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances. This approach not only corrects dataset inherent errors but also significantly improves the model’s capacity to capture the complex inter-dependencies within STG patterns.
@@ 3.3 ST-S3M: Spatial-Temporal Selective State Space Module
ST-S3M plays the significant role of adaptive STG feature selection, serving as the counterpart of Attention mechanism. The architecture of ST-S3M is shown in Figure 1. After being processed by KFGN module for dynamic spatiotemporal dependency modeling and statistical-based integration & upgrading. The generated embedding $\hat{h}_{fuse}$ is sent to ST-S3M for input-specific dynamic feature selection. In ST-S3M, the input STG embedding passes through a Linear Layer, followed by a splitting operation. Let $b$ denotes batch size, $l$ denotes sequence length, $d_{model}$ be the feature dimension, $d_{inner}$ be the model inner feature dimension, we have:
@@
$h_{\text{main}}^{{}^{\prime}}=\text{SiLU}(\text{Conv1D}(h_{\text{main}}))$ (12)
where $h_{\text{main}}^{{}^{\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to graph state space selection mechanism:
$h_{\text{sssm}}=\text{GSSSM}(h_{main}^{{}^{\prime}})$ (13)
where $h_{\text{sssm}}\in\mathbb{R}^{b\times l\times d_{inner}}$. Meanwhile, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$. The fused result is transformed through a Linear projection:
@@
The GSSSM in ST-S3M plays a main role in adaptive spatiotemporal feature selection. We detail the parameter computation and update process of the Graph State Space Selection Mechanism in Algorithm 2. Graph Selective Scan Algorithm (Algorithm 1) which receives graph information feed-forward is the most significant step in the state space selection process. We detail it in the following.
Graph Selective Scan Algorithm. The algorithm is a further extension of the basic selective scan, which integrates dynamic graph information generated by KFGN into state-space selection & update procedure, enhancing Mamba’s capability in capturing STG dependencies. The key steps and modifications are as follows:
We start with obtaining the dimension of the input tensor $u\in\mathbb{R}^{(b,l,d_{in})}$, where $b$ is the batch size, $l$ denotes the sequence length, $d_{in}$ is the input features dim; and the second dim of $\mathbb{A}$ is denoted as $n$.
[…]
Following that, we Discretize the continuous parameters $\mathbb{A}$ and $\mathbb{B}$ as:
[…]
where einsum denotes the Einstein Summation Convention. $\text{delta}\mathbb{A}$ is the updated state transition matrix, $\text{delta}\mathbb{B}_{u}$ is the updated control matrix.
Then, Iterative State Update is executed on state $x\in\mathbb{R}^{b\times d_{in}\times n}$. For time step from $i$ to $l$, we have:
$x$ $\leftarrow\text{delta}\mathbb{A}[:,i]\times x+\text{delta}\mathbb{B}_{u}[:,i]$ (16)
$z$ $\leftarrow\text{einsum}(x,\mathbb{C}[:,i,:])$
where $z$ is the current output computed by einstein summation convention. $z$ is then appended to the output list $z_{s}$. The list of outputs $z_{s}$ is stacked to form the output tensor $z$, and finally we add the direct gain $\mathbb{D}$ to the final output as: $z\leftarrow z+u\times\mathbb{D}$.
The Graph Selective Scan algorithm offers several advantages including: enhanced spatiotemporal dependency modeling, adaptability to changing graph structures, and improved forecasting accuracy, making it particularly suitable for spatiotemporal graph learning tasks.
@@ 4.1 Dataset Statistics and Baseline Methods
Datasets. We meticulously select three real-world STG datasets from California road network speed records, Hangzhou metro system entry/exit records, and weather station records across mainland China. Namely PeMS04 Song et al. (2020), HZMetro Liu et al. (2022), and KnowAir Wang et al. (2020), respectively. The detailed dataset statistics and descriptions are summarized in Table 1.
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 48 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-4}$, with a weight decay of $1e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 100. The number of GS3B Encoder Layer is set as $N=4$. We use the historical ground-truth data from the past 12 steps (for historical recent/period/trend scales) to forecast the future 12 time steps. Mean Squared Error (MSE) is employed as the Loss Function for optimization. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
4.3 Evaluation Metrics
We employ RMSE/MAE/MAPE/$std_{\text{MAE}}$ as the deterministic evaluation metrics, all the indicators are smaller the better; $\rm{\Delta_{\text{VAR}}}$ (Change in Variance Accounted For) and $\rm{R^{2}}$ (Coefficient of Determination) as statistic-based evaluation metric, where $\rm{\Delta_{\text{VAR}}}$ is the smaller the better, $\rm{R^{2}}$ is the larger the better. The reported results in Experiment section are the average value of 10 runs. The detailed indicator calculation methods are provided in Appendix B because of limited space.